Calculate the (a) resistivity and the (b) conductivity of a material that has a length of 15 m and a cross-sectional area of 5 x 10 m² and a resistance (R) of 5 ohms


1
Expert's answer
2021-05-17T18:57:44-0400

Solution:-

Gives

Length (l)=15meter

Cross section area (A)="5\times10 m^2=50m^2"

R=5 ohms

Part(a)

Resistivity

"\rho=\frac{RA}{l}\rightarrow(1)"

Put value in eqution (1)


"\rho=\frac{{5}\times{50}}{15}\Omega m"


"\rho=16.66\Omega m"

Part (b)

Conductivity

"\sigma=\frac{1}{\rho}"

"\sigma=\frac{15}{5\times50} mho/meter"

"\sigma=0.06mho/meter"